A card is selected from a deck of 52 playing cards. Find the probability of selecting
a) a seven given the card is a not a face card. (An ace is not a face card.)
b) a spade given the card is red.



Answer by checkley77(12844)

4/(52-16)=4/36=1/9=.1111=11.11% chance of selecting a 7.
26/0 not possible.